What is the width of a strip that has been mowed around a rectangulare field that is 60 ft. by 80 ft. if half the lawn has not yet been mowed?
Answer by Emsley(35)

I f thr strip is of width x ft then the inside rectangle has length (80-2x) and width ((60-2x) But its area = 2400
So (80-2x)(60-2x)=2400 This tidies up to 4x^2-280x+2400=0
Divide both sides by 4 getting x^2-70x+600=0 Factorise (x-10)(x-60)=0
x= 10 or 60 But x cannot be 60 so x=10